### 10 Mar 2026 v1

# ACS CONDITION ON MINIMAL ISOPARAMETRIC HYPERSURFACES OF POSITIVE RICCI CURVATURE IN UNIT SPHERES

Abstract.
Motivated by the Schoen–Marques–Neves conjecture relating the Morse index of a minimal hypersurface to its first Betti number in ambient manifolds with positive Ricci curvature, we study the Ambrozio–Carlotto–Sharp (ACS) criterion for index growth. We verify a sufficient pointwise ACS inequality for several families of minimal isoparametric hypersurfaces with positive Ricci curvature in the unit sphere. As a consequence, for any closed embedded minimal hypersurface $M^{n}$ in such an ambient manifold $\mathcal{N}^{n+1}\subset\mathbb{S}^{n+2}\subset\mathbb{R}^{n+3}$, one has $\operatorname{index}(M)\geq\frac{2}{d(d-1)}\,b_{1}(M),d=n+3,$ where $b_{1}(M)$ is the first Betti number of $M$.
Key words and phrases:
Differential geometry, isoparametric theory, Morse index, minimal hypersurface
2020 Mathematics Subject Classification:
53C42
1. Introduction
In 1970, Lawson [11] constructed minimal surfaces in $S^{3}$ of arbitrary genus. This work initiated a new direction in differential geometry by moving the study of minimal surfaces from Euclidean space to positively curved ambient spaces. Later, Hsiang and Lawson [7] extended related techniques to higher-dimensional spheres. On the other hand, Lawson’s work also clarified topological restrictions for minimal surfaces in positively curved manifolds and influenced the study of incompressible surfaces in three-manifolds by Schoen and Yau [18].
Besides the existence problem, an equally important question is to understand the geometric and topological properties of minimal hypersurfaces. Among these, the Morse index is a central analytical invariant: it measures stability and reflects deep information about the geometry of the hypersurface. It has been conjectured that the index controls the basic topology of minimal hypersurfaces in an effective way.
Conjecture (Schoen-Marques-Neves, see [12]).
Let $(N^{n+1},g)$ be a closed Riemannian manifold with positive Ricci curvature. Then there exists a positive constant $C$ such that, for every closed embedded minimal hypersurface $M^{n}\subset N^{n+1}$,
$\operatorname{index}(M)\geq C\,b_{1}(M).$
A particular ambient manifold for which this conjecture is known is the round sphere. This was proved by Savo [17], building on earlier work of Ros [16]. For flat three-dimensional tori, Ros proved that the index of a closed minimal surface is bounded from below by an affine function of the genus. Since, for a closed orientable surface, the first Betti number is twice the genus, this is a close analogue of the conjectural picture above.
Ambrozio, Carlotto and Sharp [1] developed a general framework to study this conjecture in other positively curved ambient manifolds. Their method gives a lower bound for the Morse index in terms of the first Betti number under a suitable extrinsic inequality.
Theorem 1.1 (Ambrozio-Carlotto-Sharp [1]).
Let $(N^{n+1},g)$ be a Riemannian manifold isometrically embedded in some Euclidean space $\mathbb{R}^{d}$. Let $M^{n}$ be a closed embedded minimal hypersurface of $N^{n+1}$. Assume that for every nonzero vector field $X$ on $M$,
(1) $\int_{M}\Bigg[\sum_{k=1}^{n}R^{N}(e_{k},X,e_{k},X)+\operatorname{Ric}_{N}(\nu,\nu)|X|^{2}\Bigg]\,dM$
$>$ $\int_{M}\Bigg[\sum_{k=1}^{n}|\Pi(e_{k},X)|^{2}+\sum_{k=1}^{n}|\Pi(e_{k},\nu)|^{2}|X|^{2}\Bigg]\,dM.$
where $R^{N}$ is the curvature tensor of $N$, $\Pi$ is the second fundamental form of the embedding $N\hookrightarrow\mathbb{R}^{d}$, $\nu$ is a local unit normal vector field along $M$, and $\{e_{k}\}_{k=1}^{n}$ is a local orthonormal frame on $M$. Then
$\operatorname{index}(M)\geq\frac{2}{d(d-1)}\,b_{1}(M).$
For convenience, we refer to (1) as the ACS condition. The same strategy has since been extended to free boundary and weighted settings; see [2, 4, 8, 9].
In this paper we apply the ACS framework to minimal isoparametric hypersurfaces with positive Ricci curvature in the unit sphere. We hope to provide further evidence for the Schoen-Marques-Neves conjecture by adding new examples in this class.
For an isoparametric hypersurface $N^{n+1}\subset S^{n+2}$, the number $g$ of distinct principal curvatures can only be $1,2,3,4$ or $6$, and the multiplicities satisfy $m_{i}=m_{i+2}$ (indices modulo $g$). Hence there are at most two distinct multiplicities, say $m_{1}$ and $m_{2}$; see [3, 10]. For a minimal isoparametric hypersurface $N^{n+1}\subset S^{n+2}$, the cases with positive Ricci curvature are the following [15]:
  1. (1)
    when $g=1$, the Ricci curvature is obviously positive;
  2. (2)
    when $g=2$, the Ricci curvature is positive unless $N$ is $S^{1}(r_{1})\times S^{n}(r_{2})$;
  3. (3)
    when $g=3$, the multiplicities are equal and the Ricci curvature is positive for $m_{1}=m_{2}>1$;
  4. (4)
    when $g=4$, the Ricci curvature is positive provided $m_{1},m_{2}\geq 2$.
Our main result is the following.
Theorem 1.2.
Let $N^{n+1}\subset S^{n+2}$ be a minimal isoparametric hypersurface. Then the following statements hold.
  1. (1)
    If $g=3$ and $m_{1}=m_{2}=4$ or $8$, then $N$ satisfies the ACS condition.
  2. (2)
    If $g=4$ and $m_{1},m_{2}\geq 5$, then $N$ satisfies the ACS condition.
Consequently, if $M^{n}$ is a closed minimal hypersurface of any of the above ambient manifolds, then
$\operatorname{index}(M)\geq\frac{2}{d(d-1)}\,b_{1}(M),$
where $d=n+3$ for the natural embedding $N\subset S^{n+2}\subset\mathbb{R}^{n+3}$.
Gorodski, Mendes and Radeschi [6] also produced many examples supporting the Schoen-Marques-Neves conjecture. They obtained robust index lower bounds for minimal hypersurfaces in several isoparametric and symmetric ambient spaces by a different method. Our contribution is instead to verify a sufficient ACS inequality for the ambient manifold itself.Our arguments in the $g=4$ case are deliberately coarse: we impose a uniform bound on the largest absolute principal curvature. This yields a clean sufficient condition $\min\{m_{1},m_{2}\}\geq 5$.
The cases not covered by Theorem 1.2 remain open from the point of view of the present computation. More precisely:
  1. (1)
    for $g=3$, the case $m_{1}=m_{2}=2$ is not decided by our argument, while for $m_{1}=m_{2}=1$ the Ricci curvature is not positive;
  2. (2)
    for $g=4$, our estimate does not decide the ACS condition when at least one multiplicity lies in $\{2,3,4\}$;
  3. (3)
    for $g=6$, the Ricci curvature is not positive.
2. Preliminaries
An isoparametric hypersurface is a hypersurface that arises as a regular level set of an isoparametric function. More precisely, a hypersurface $N^{n+1}$ in a Riemannian manifold $\widetilde{N}^{n+2}$ is called isoparametric if there exists a nonconstant smooth function $f$ on $\widetilde{N}$ and smooth functions $a,b$ such that
  1. (1)
    $N=f^{-1}(c)$ for some regular value $c$ of $f$;
  2. (2)
    $|\nabla f|^{2}=b(f)$;
  3. (3)
    $\Delta f=a(f)$.
In a space form, this is equivalent to requiring that all principal curvatures are constant.
For isoparametric hypersurfaces in the unit sphere $S^{n+2}$, the possible numbers of distinct principal curvatures are $1,2,3,4$ and $6$. The cases may be summarized as follows (see [5, 3, 13, 14, 10]):
  1. (1)
    if $g=1$, then $N$ is totally umbilic;
  2. (2)
    if $g=2$, then $N$ is a Clifford hypersurface $S^{k}(r)\times S^{n+1-k}(\sqrt{1-r^{2}})$;
  3. (3)
    if $g=3$, then all multiplicities are equal and can be $1,2,4$ or $8$;
  4. (4)
    if $g=6$, then all multiplicities are equal and can be $1$ or $2$;
  5. (5)
    if $g=4$, then the examples are either homogeneous or of OT-FKM type; among the multiplicity pairs, only $(2,2)$ and $(4,5)$ are not of OT-FKM type.
For OT-FKM hypersurfaces, the multiplicities are of the form $(m_{1},m_{2})=(m_{1},\ell-m_{1}-1)$ with $\ell\in\mathbb{N}^{*}$.
We will also use the following result of Münzner.
Theorem 2.1 (Münzner, see [5, 13, 14]).
Let $M^{n}$ be an isoparametric hypersurface with $g$ distinct principal curvatures in $S^{n+1}$. Let
$\lambda_{1}=\cot\theta_{1}>\cdots>\lambda_{g}=\cot\theta_{g},\qquad\theta_{i}\in(0,\pi),$
be the principal curvatures, with multiplicities $m_{1},\dots,m_{g}$. Then:
  1. (1)
    there exists a homogeneous polynomial $F$ of degree $g$ on $\mathbb{R}^{n+2}$, called the Cartan-Münzner polynomial, such that
    $|\nabla F|^{2}=g^{2}r^{2g-2},\qquad\Delta F=\frac{m_{2}-m_{1}}{2}\,g^{2}r^{g-2},$
    where $r$ is the radial function on $\mathbb{R}^{n+2}$;
  2. (2)
    if $f=F|_{S^{n+1}}$, then the range of $f$ is [1,1][-1,1], the only critical values are $\pm 1$, and the critical sets $M_{\pm}=f^{-1}(\pm 1)$ are connected focal submanifolds;
  3. (3)
    for each $c\in(-1,1)$, the level set $M_{c}=f^{-1}(c)$ is an isoparametric hypersurface, the multiplicities satisfy $m_{i}=m_{i+2}$ (indices modulo $g$), and
    $\theta_{i}=\theta_{1}+\frac{i-1}{g}\pi;$
    in particular, if $g$ is odd then all multiplicities are equal;
  4. (4)
    the unique minimal member in the isoparametric family is $M_{c_{0}}=f^{-1}(c_{0})$, where
    $c_{0}=\frac{m_{2}-m_{1}}{m_{1}+m_{2}}.$
3. The computations
Let $N^{n+1}\subset S^{n+2}$ be a minimal isoparametric hypersurface. To check the ACS condition, it is enough to study the pointwise integrand in (1).
Fix a closed minimal hypersurface $M^{n}\subset N^{n+1}$. Let $\nu$ be a local unit normal vector field of $M$ in $N$, and let $\{e_{k}\}_{k=1}^{n}$ be a local orthonormal frame on $M$. We define
(2) $\Delta(X,\nu)$={} $\sum_{k=1}^{n}R^{N}(e_{k},X,e_{k},X)+\operatorname{Ric}_{N}(\nu,\nu)|X|^{2}$
$-\sum_{k=1}^{n}|\Pi(e_{k},X)|^{2}-\sum_{k=1}^{n}|\Pi(e_{k},\nu)|^{2}|X|^{2}.$
If $\Delta(X,\nu)>0$ for every nonzero tangent vector $X$, then $N$ satisfies the ACS condition.
Let $\{\varepsilon_{i}\}_{i=1}^{n+1}$ be a local orthonormal frame of principal directions of $N\subset S^{n+2}$, with corresponding principal curvatures $\{\lambda_{i}\}_{i=1}^{n+1}$. Write
$X=\sum_{i=1}^{n+1}x_{i}\varepsilon_{i},\qquad\nu=\sum_{i=1}^{n+1}\nu_{i}\varepsilon_{i}.$
Since (2) is homogeneous in $X$, we may assume $|X|=1$. We also have $|\nu|=1$ and $\langle X,\nu\rangle=0$.
Proposition 3.1.
With the notation above,
$\Delta(X,\nu)={}$ $(2n-2)-2\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}^{2}-2\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}^{2}+2\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}$
$\qquad+\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)^{2}-\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)\Big(\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}\Big).$
Proof.
Let $\eta$ be a local unit normal vector field of $N$ in $S^{n+2}$, and let $\xi$ be the unit normal vector field of $S^{n+2}$ in $\mathbb{R}^{n+3}$. Then, for tangent vectors $U,V\in TN$,
$\Pi(U,V)=\Pi^{\eta}(U,V)+\Pi^{\xi}(U,V)=\langle AU,V\rangle\eta+\langle U,V\rangle\xi,$
where $A$ is the shape operator of $N\subset S^{n+2}$. In the principal frame,
$A(\varepsilon_{i})=\lambda_{i}\varepsilon_{i},\qquad\Pi(\varepsilon_{i},\varepsilon_{j})=\lambda_{i}\delta_{ij}\eta+\delta_{ij}\xi.$
We first compute the curvature term. Since $\{e_{1},\dots,e_{n},\nu\}$ is an orthonormal frame of $TN$,
$\sum_{k=1}^{n}R^{N}(e_{k},X,e_{k},X)=\operatorname{Ric}_{N}(X,X)-R^{N}(\nu,X,\nu,X).$
By the Gauss equation for $N\subset S^{n+2}$,
$K(\varepsilon_{i},\varepsilon_{j})=1+\lambda_{i}\lambda_{j}\qquad(i\neq j),$
and the minimality of $N$ implies $\sum_{i}\lambda_{i}=0$. Therefore,
$\operatorname{Ric}_{N}(\varepsilon_{i},\varepsilon_{i})$ $=\sum_{j\neq i}K(\varepsilon_{j},\varepsilon_{i})=\sum_{j\neq i}(1+\lambda_{j}\lambda_{i})=n+\lambda_{i}\sum_{j\neq i}\lambda_{j}=n-\lambda_{i}^{2}.$
Hence
(3) $\operatorname{Ric}_{N}(X,X)=\sum_{i=1}^{n+1}(n-\lambda_{i}^{2})x_{i}^{2}.$
On the other hand,
$\Pi(\nu,\nu)$ $=\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)\eta+\xi,$
$\Pi(X,X)$ $=\Big(\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}\Big)\eta+\xi,$
$\Pi(\nu,X)$ $=\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)\eta,$
since $\langle\nu,X\rangle=0$. Applying the Gauss equation again, we obtain
$R^{N}(\nu,X,\nu,X)$ $=\langle\Pi(\nu,\nu),\Pi(X,X)\rangle-|\Pi(\nu,X)|^{2}$
(4) $=\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)\Big(\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}\Big)+1-\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}.$
Combining (3) and (4), we get
$\sum_{k=1}^{n}R^{N}(e_{k},X,e_{k},X)={}$ $\sum_{i=1}^{n+1}(n-\lambda_{i}^{2})x_{i}^{2}-\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)\Big(\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}\Big)$
(5) $\qquad-1+\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}.$
Moreover,
(6) $\operatorname{Ric}_{N}(\nu,\nu)|X|^{2}=\sum_{i=1}^{n+1}(n-\lambda_{i}^{2})\nu_{i}^{2}.$
We next compute the second fundamental form terms. Since
$\{e_{1},\dots,e_{n},\nu\}$ is an orthonormal basis of $TN$,
$\sum_{k=1}^{n}|\Pi(e_{k},X)|^{2}$ $=\sum_{i=1}^{n+1}|\Pi(\varepsilon_{i},X)|^{2}-|\Pi(\nu,X)|^{2},$
$\sum_{k=1}^{n}|\Pi(e_{k},\nu)|^{2}$ $=\sum_{i=1}^{n+1}|\Pi(\varepsilon_{i},\nu)|^{2}-|\Pi(\nu,\nu)|^{2}.$
Now,
$\Pi(\varepsilon_{i},X)=x_{i}\lambda_{i}\eta+x_{i}\xi,\qquad\Pi(\varepsilon_{i},\nu)=\nu_{i}\lambda_{i}\eta+\nu_{i}\xi,$
so that
$|\Pi(\varepsilon_{i},X)|^{2}=x_{i}^{2}(\lambda_{i}^{2}+1),\qquad|\Pi(\varepsilon_{i},\nu)|^{2}=\nu_{i}^{2}(\lambda_{i}^{2}+1).$
Therefore,
$\sum_{k=1}^{n}|\Pi(e_{k},X)|^{2}+\sum_{k=1}^{n}|\Pi(e_{k},\nu)|^{2}|X|^{2}={}$ $\sum_{i=1}^{n+1}x_{i}^{2}(\lambda_{i}^{2}+1)+\sum_{i=1}^{n+1}\nu_{i}^{2}(\lambda_{i}^{2}+1)$
$-\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}$
(7) $-\Bigg[\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)^{2}+1\Bigg].$
Finally, substituting (5), (6) and (7) into (2), and using $\sum_{i}x_{i}^{2}=\sum_{i}\nu_{i}^{2}=1$, we obtain the claimed formula. ∎
Proof of Theorem 1.2.
For simplicity, write
$\Delta:={}$ $(2n-2)-2\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}^{2}-2\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}^{2}+2\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}$
$\qquad+\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)^{2}-\Big(\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}\Big)\Big(\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}\Big).$
By Proposition 3.1, it suffices to show that $\Delta>0$.
(1) The case $g=3$. For a minimal isoparametric hypersurface with $g=3$, the principal curvatures are
$\lambda_{1}=\sqrt{3},\qquad\lambda_{2}=0,\qquad\lambda_{3}=-\sqrt{3}.$
If the common multiplicity is $m=4$ or $8$, then
$n=3m-1\geq 11,\qquad 2n-2\geq 20.$
Since $|X|=|\nu|=1$ and $\langle X,\nu\rangle=0$, we have
$\sum_{i}x_{i}^{2}=\sum_{i}\nu_{i}^{2}=1,\qquad\sum_{i}x_{i}\nu_{i}=0.$
Hence
$\sum_{i}x_{i}^{2}\lambda_{i}^{2}\leq 3,\qquad\sum_{i}\nu_{i}^{2}\lambda_{i}^{2}\leq 3,\qquad\left|\Big(\sum_{i}\nu_{i}^{2}\lambda_{i}\Big)\Big(\sum_{i}x_{i}^{2}\lambda_{i}\Big)\right|\leq 3.$
Discarding the nonnegative terms in the expression for $\Delta$, we obtain
$\Delta\geq(2n-2)-6-6-3\geq 20-6-6-3>0.$
Therefore $N$ satisfies the ACS condition when $g=3$ and $m=4$ or $8$.
(2) The case $g=4$. Set
$s:=m_{1}+m_{2},\qquad n=2s-1.$
By Theorem 2.1, the principal curvatures are
$\lambda_{i}=\cot\Big(\theta_{1}+\frac{i-1}{4}\pi\Big),\qquad i=1,2,3,4,$
where, for the minimal leaf,
$\cos 4\theta_{1}=c_{0}=\frac{m_{2}-m_{1}}{m_{1}+m_{2}}.$
Let
$a^{2}:=\max_{1\leq i\leq 4}\lambda_{i}^{2}=\max\{\lambda_{1}^{2},\lambda_{4}^{2}\}.$
Since
$\sum_{i}x_{i}^{2}\lambda_{i}^{2}\leq a^{2},\qquad\sum_{i}\nu_{i}^{2}\lambda_{i}^{2}\leq a^{2},\qquad\left|\sum_{i}x_{i}^{2}\lambda_{i}\right|\leq a,\qquad\left|\sum_{i}\nu_{i}^{2}\lambda_{i}\right|\leq a,$
Proposition 3.1 gives
(8) $\Delta\geq(2n-2)-5a^{2}.$
Thus it is enough to require
(9) $a^{2}<\frac{2n-2}{5}=\frac{4}{5}(s-1).$
We do not assume any ordering between $m_{1}$ and $m_{2}$. Since $0<2\theta_{1}<\pi/2$, we have
$\cos 2\theta_{1}=\sqrt{\frac{1+\cos 4\theta_{1}}{2}}=\sqrt{\frac{m_{2}}{s}},\qquad\sin 2\theta_{1}=\sqrt{\frac{1-\cos 4\theta_{1}}{2}}=\sqrt{\frac{m_{1}}{s}}.$
Therefore,
$\lambda_{1}^{2}$ $=\cot^{2}\theta_{1}=\frac{1+\cos 2\theta_{1}}{1-\cos 2\theta_{1}}=\frac{1+\sqrt{m_{2}/s}}{1-\sqrt{m_{2}/s}},$
$\lambda_{4}^{2}$ $=\cot^{2}\Big(\theta_{1}+\frac{3\pi}{4}\Big)=\frac{1+\sin 2\theta_{1}}{1-\sin 2\theta_{1}}=\frac{1+\sqrt{m_{1}/s}}{1-\sqrt{m_{1}/s}}.$
If we set
$p:=\max\{m_{1},m_{2}\},\qquad q:=\min\{m_{1},m_{2}\}=s-p,$
then
$a^{2}=\frac{1+\sqrt{p/s}}{1-\sqrt{p/s}}.$
Condition (9) is therefore equivalent to
$\sqrt{\frac{p}{s}}<\frac{\frac{4}{5}(s-1)-1}{\frac{4}{5}(s-1)+1},$
and hence to
(10) $q=s-p>s\left[1-\left(\frac{\frac{4}{5}(s-1)-1}{\frac{4}{5}(s-1)+1}\right)^{2}\right]=\frac{80s(s-1)}{(4s+1)^{2}}.$
Now,
$5(4s+1)^{2}-80s(s-1)=120s+5>0,$
so
$\frac{80s(s-1)}{(4s+1)^{2}}<5.$
Thus (10) is certainly satisfied if $q\geq 5$, i.e. if
$\min\{m_{1},m_{2}\}\geq 5.$
Equivalently, if $m_{1},m_{2}\geq 5$, then (9) holds, and therefore (8) yields $\Delta>0$. This proves the ACS condition for the case $g=4$ with $m_{1},m_{2}\geq 5$.
The index estimate then follows immediately from Theorem 1.1. ∎
Remark 3.2.
The argument above leaves the following cases unresolved by this method: the case $g=3$ with multiplicity $m=2$, and the case $g=4$ when at least one multiplicity belongs to $\{2,3,4\}$. The cases $g=3$ with $m=1$ and $g=6$ are excluded because the ambient Ricci curvature is not positive.